# RALLY: Role-Adaptive LLM-Driven Yoked Navigation for Agentic UAV Swarms

### v1

@@ Abstract
Intelligent control of Unmanned Aerial Vehicles (UAVs) swarms has emerged as a critical research focus, and it typically requires the swarm to navigate effectively while avoiding obstacles and achieving continuous coverage over multiple mission targets. Although traditional Multi-Agent Reinforcement Learning (MARL) approaches offer dynamic adaptability, they are hindered by the semantic gap in numerical communication and the rigidity of homogeneous role structures, resulting in poor generalization and limited task scalability. Recent advances in Large Language Model (LLM)-based control frameworks demonstrate strong semantic reasoning capabilities by leveraging extensive prior knowledge. However, due to the lack of online learning and over-reliance on static priors, these works often struggle with effective exploration, leading to reduced individual potential and overall system performance. To address these limitations, we propose a Role-Adaptive LLM-Driven Yoked navigation algorithm RALLY. Specifically, we first develop an LLM-driven semantic decision framework that uses structured natural language for efficient semantic communication and collaborative reasoning. Afterward, we introduce a dynamic role-heterogeneity mechanism for adaptive role switching and personalized decision-making. Furthermore, we propose a Role-value Mixing Network (RMIX)-based assignment strategy that integrates LLM offline priors with MARL online policies to enable semi-offline training of role selection strategies. Experiments in the Multi-Agent Particle Environment (MPE) environment and a Software-In-The-Loop (SITL) platform demonstrate that RALLY outperforms conventional approaches in terms of task coverage, convergence speed, and generalization, highlighting its strong potential for collaborative navigation in agentic multi-UAV systems.
Index Terms:
@@ I Introduction
Nowadays, Unmanned Aerial Vehicles (UAVs) have demonstrated significant potential in multi-agent pursuit-evasion game (e.g., disaster responses [1]). Typically, the UAV swarm shall have the capability to avoid pursuing enemies and environmental obstacles while moving towards multiple target areas in certain formations, which is often known as a Dynamic Swarm coordination with Cooperative Evasion and Formation Coverage (DS-CEFC) task [2]. Nevertheless, the underlying difficulty in coordinating roles and decision-making across multiple agents in swarms poses significant challenges. For example, traditional control-based algorithms [3, 4] are contingent on fully connected topologies and lack the required adaptability and scalability to large-scale dynamic environments, thus degrading the practicality in the real world [5]. Meanwhile, despite the incorporation of inter-agent communications [6, 7] and cooperation [8, 9], it still remains inevitable for decentralized Multi-Agent Reinforcement Learning (MARL) to yield conflicting roles and decisions [10, 11]. Given the inherent semantic reasoning capabilities and pretrained experience of LLM [12, 13], they offer a promising alternative to mitigate the critical issues of conflicting roles and inconsistent decision-making that plague purely MARL-based approaches.
I-A Related Works
I-A1 MARL-Based UAV Swarm Control
@@
Deep Reinforcement Learning (DRL)[14] has significantly enhanced agent adaptability in complex tasks. However, in Multi-Agent Systems (MAS), critical challenges, such as environmental non-stationarity, rapidly expanding state spaces, and difficulty in credit assignment, hinder the ability of traditional methods to learn effective cooperative policies. To address this problem, the Centralized Training with Decentralized Execution (CTDE) paradigm is introduced with exemplary algorithms like MADDPG[15] and MAPPO[16]. During training, it also involves techniques such as policy distillation[17] and imitation learning[18] to improve coordination in complex obstacle-laden environments. Nevertheless, these approaches struggle to assess an individual agent’s contribution, as they typically optimize a global value function while neglecting the importance of localized utility. Value decomposition methods (e.g., VDN[8], QMIX[19], QTRAN[20]) ameliorate this issue by decomposing the joint value function, thereby enabling analysis of individuals’ contribution to cooperative decision-making. Furthermore, more advanced attention mechanisms or neural communication protocols (e.g., TarMAC[9], IMANet[21], DAACMP[22]) can boost the effectiveness of filter messages. However, the direct communication of numerical vectors[9, 21, 22], which lacks interpretability and cannot convey task semantics, leads to information redundancy and bandwidth bottlenecks [6, 7, 23, 24], and greatly limits algorithm generalization. Although the leader–follower architecture[25] enables role differentiation, its static role assignments lack the flexibility to adapt to dynamic environmental conditions and varying formation sizes. Therefore, researchers resort to integrating hierarchical control with consensus inference[26, 27, 2], so as to simplify the inference interpretability and boost convergence speed. However, agents in these methods commonly remain homogeneous, and the policy networks learned via CTDE cannot inherently leverage the advantages of the natural heterogeneity in swarms. Therefore, building a scalable, efficient, dynamically adaptive, and interpretable heterogeneous multi-agent collaboration mechanism for UAV swarm control remains an open problem.
I-A2 LLM-Based Multi-Agent Systems
@@
Large Language Models (LLMs) have demonstrated near-human performance[12] in complex reasoning and planning tasks, providing new impetus for environment understanding and decision explainability in UAV swarms[13]. Leveraging vast prior knowledge, LLMs can not only be used for single-agent path planning (e.g., CoNavGPT[28], RoCo[29]), but also facilitate high-level communication with low-level trajectory planning, significantly improving task efficiency, adaptability, and generalization[30, 31, 32]. Moreover, in MetaGPT[33], CAMEL[34] and ChatDev[35], LLMs can decompose complex missions into a number of subtasks handled collaboratively by different agents, thereby reducing “hallucinations” [36] and enhancing the ability to solve complex problems. More importantly, these LLM-driven Multi-Agent (LLM-MA) systems[37, 38, 39] often customize LLMs into specialized or personalized roles. Therefore, through multi-agent collaboration, they replicate human-like collective intelligence and further enhance overall situational understanding and decision explainability, making inter-agent interactions more meaningful. However, LLM-based decision-making still heavily relies on prior knowledge and lacks exploration, often getting stuck in local optima[40]. Additionally, these methods leverage individual memory and self-evolution mechanisms to optimize agents independently, while neglecting potential synergistic effects arising from multi-agent interactions. Furthermore, applying LLM-MA to build global consensus for DS-CEFC remains scarce, though some studies[39] start to focus on consensus negotiation among agents in a simplified scenario. On the other hand, some approaches[37] utilize diverse “personalities”, empowered by the LLM, for creative collaboration, nevertheless the underlying fixed definitions of roles still struggle to accommodate dynamic task switching in DS-CEFC. In other words, enabling dynamic role adaptation with balanced exploration and exploitation remains a critical challenge for DS-CEFC.
I-A3 Integration of MARL and LLM
@@
There has shed some light on deeply integrating MARL and LLMs[41]. Prominently, semantic capabilities in LLMs can be leveraged as natural language interfaces to bridge human feedback and MAS [42], but the consensus inference therein often relies more on human supervision than on autonomous collaboration. Other studies distill LLM knowledge into smaller executable models or in combination with human-in-the-loop feedback, using LLMs for policy generation to accelerate MARL training and improve performance in complex environments [43, 44, 45]. Nevertheless, these methods often depend on offline human annotations and feedback, making them ill-suited for dynamic real-time changes, meanwhile agent roles remain homogeneous, with only the number of agents being scaled up. Some recent work treats LLMs as the core of heterogeneous agents in MARL[12], assigning different “personalities” to agents. But these roles tend to be fixed and cannot adapt to environmental changes in DS-CEFC. Additionally, deploying LLMs in UAV swarms significantly strains computational resources and energy, limiting their practical use [46]. Thus, it is meaningful to calibrate superior means to integrate MARL and LLMs, thus better harnessing heterogeneous swarm intelligence for improved generalization and adaptability in DS-CEFC.
I-B Contributions
@@
To accomplish consensus inference of multiple UAVs with heterogeneous roles for DS-CEFC, we propose a novel LLM-MARL-integrated framework called RALLY (Role-Adaptive LLM-Driven Yoked navigation). Built on our previous work[2], which unifies target selection, obstacle-avoidance navigation, and flight-control execution, RALLY serves as an significantly enhanced high-level consensus inference module towards establishing role-adaptive cooperative navigation. Specifically, RALLY comprises two core modules: an LLM-based two-stage semantic reasoning module and a Role-value Mixing Network (RMIX)-based credit-distribution mechanism. The integration of semantic reasoning in LLM and policy learning in MARL makes it qualified for coordinating roles and decision-making across swarms for DS-CEFC. The main contributions of this work can be summarized as follows:
@@
  • To meet the distributed deployment and parallel inference demands on edge devices, we propose a capacity-migration algorithm that significantly reduces runtime memory footprint. We validate RALLY in the Multi-Particle Environment (MPE)[47] and the Software-In-The-Loop (SITL) platform[48] based on Gazebo-ROS-PX4. Experimental results demonstrate that RALLY outperforms latest MARL[2] and pure LLM-driven approaches[28, 49] in terms of task completion, convergence speed, generalization, and interpretability.
I-C Paper Structure
The remainder of this paper is organized as follows. Section II introduces the system model and formulates the problem. Section III outlines the algorithmic framework of RALLY by elaborating on the RMIX role assignment mechanism and context-migration algorithm. Section IV presents experimental results and discussion. Finally, Section V summarizes this work with possible research directions.
II System Model and Problem Formulation
II-A System Model
Beforehand, commonly used variables are summarized in Table I.
TABLE I: Mainly used notations in this paper.
@@ II-A System Model
$\bm{p}_{e}^{t}$, $\bm{v}_{e}^{t}$ Anemy’s position/velocity
@@
$\hat{g}_{i}^{t}$ Initial goal intention of agent $i$
@@
$\text{LLM}_{\text{HC}}$ Two Stege LLM-based consensus inference
@@
where $\bm{p}_{i}^{t}=[p_{x_{i}}^{t},p_{y_{i}}^{t}]$ and $\bm{v}_{i}^{t}=[v_{x_{i}}^{t},v_{y_{i}}^{t}]$ are the 2D position and velocity of agent $i$, respectively, while $\bm{p}_{e}^{t}=[p_{x_{e}}^{t},p_{y_{e}}^{t}]$, $\bm{v}_{e}^{t}=[\bm{v}_{x_{e}}^{t},\bm{v}_{y_{e}}^{t}]$ are those of enemy agent. Based on its local observation $\bm{o}_{i}^{t}$, agent $i$ first generates a target intention $\hat{g}_{i}^{t}\in\mathcal{T}$ and role assignment $k_{i}^{t}\in\mathcal{K}=\{\texttt{Commander},\texttt{Coordinator},\texttt{Executor}\}$ (i.e., $\bm{a}_{i}^{t}=(\hat{g}_{i}^{t},k_{i}^{t})$), and then exchanges these proposals and observation (denoted as $\bm{M}_{\text{i}}^{t}=(\bm{a}_{i}^{t},\bm{o_{i}^{t}})$) within its neighborhood set $\mathcal{N}_{i}^{t}\in\mathcal{N}$, ultimately forming a collective consensus $g_{i}^{t}\in\mathcal{T}$ through semantic negotiation. Notably, Commander focuses on maximizing individual rewards through independent decision-making, tending to select points yielding the highest personal return; Coordinator balances team and individual gains, giving priority to the Commander’s choices whenever necessary; Executor primarily adheres to the Coordinator’s guidance, reverting to its own strategy if necessary for task reliability. As discussed later, the heterogeneous role assignment contributes to the consensus inference of large-scale swarms.
@@
As shown in Fig. 2, we further adopt a hierarchical policy structure. The mid-layer policy $\bm{\pi}_{M}$ guides the obstacle avoidance and formation $\mathcal{F}_{c}$, consisting of different numbers $c\in[C_{\min},C_{\max}]$ via MARL-based navigation, while low-layer standard PID [51] flight control $\bm{\pi}_{L}$ to stabilize UAV dynamics. In this paper, we assume the availability of $\bm{\pi}_{M}$ and $\bm{\pi}_{L}$ [2], and focus on the learning of the high-level policy $\bm{\pi}_{H}$ only, which is responsible for the inferred consensus $\mathcal{G}^{t}:=\{g_{i}^{t},\forall i\in\mathcal{N}\}$.
Figure 3: Flowchart of the RALLY algorithm. (a) Taking formation $\mathcal{F}_{4}$ as an example, UAVs 1, 3, 4, and 7 adaptively form the team. (b) Based on local observations $\bm{o}_{i}^{t}$, each agent selects target intention $\hat{g}_{i}^{t}$ and role $k_{i}$. Based on background prompt descriptions $\text{X}_{\text{task}}$, reasoning logic $\text{M}_{\text{CoT}}$, and local observation prompts $\text{Y}_{\text{init}}(\bm{o}_{i}^{t})$, UAVs generate target intentions via policy $\text{LLM}_{\text{init}}$, and then determine roles $k_{i}$ through a role selection policy $\text{RL}_{\text{HI}}$. (c) After neighborhood communications, $\text{LLM}_{\text{cons}}$ refines the goal to reach a final, converged consensus decision $g_{i}^{t}$ (e.g., Region A), incorporating individual role preferences. (d) The resulting formation $\mathcal{F}_{4}$ then collectively navigates toward Region A.
II-B Problem Formulation
We define DS-CEFC task as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP) extended to a multi-agent setting with heterogeneous roles and hierarchical semantics. The joint system state at time $t$ is

### v2

@@ Abstract
Intelligent control of Unmanned Aerial Vehicles (UAVs) swarms has emerged as a critical research focus, and it typically requires the swarm to navigate effectively while avoiding obstacles and achieving continuous coverage over multiple mission targets. Although traditional Multi-Agent Reinforcement Learning (MARL) approaches offer dynamic adaptability, they are hindered by the semantic gap in black-boxed communication and the rigidity of homogeneous role structures, resulting in poor generalization and limited task scalability. Recent advances in Large Language Model (LLM)-based control frameworks demonstrate strong semantic reasoning capabilities by leveraging extensive prior knowledge. However, due to the lack of online learning and over-reliance on static priors, these works often struggle with effective exploration, leading to reduced individual potential and overall system performance. To address these limitations, we propose a Role-Adaptive LLM-Driven Yoked navigation algorithm RALLY. Specifically, we first develop an LLM-driven semantic decision framework that uses structured natural language for efficient semantic communication and collaborative reasoning. Afterward, we introduce a dynamic role-heterogeneity mechanism for adaptive role switching and personalized decision-making. Furthermore, we propose a Role-value Mixing Network (RMIX)-based assignment strategy that integrates LLM offline priors with MARL online policies to enable offline training of role selection strategies. Experiments in the Multi-Agent Particle Environment (MPE) and a Software-In-The-Loop (SITL) platform demonstrate that RALLY outperforms conventional approaches in terms of task coverage, convergence speed, and generalization, highlighting its strong potential for collaborative navigation in agentic multi-UAV systems.
{IEEEkeywords}
@@ 1 Introduction
Nowadays, Unmanned Aerial Vehicles (UAVs) have demonstrated significant potential in multi-agent pursuit-evasion game (e.g., disaster responses [1]). Typically, the UAV swarm shall have the capability to avoid pursuing enemies and environmental obstacles while moving towards multiple target areas in certain formations, which is often known as a Dynamic Swarm coordination with Cooperative Evasion and Formation Coverage (DS-CEFC) task [2]. By enabling UAVs to adjust their functions in response to real-time environmental conditions, dynamic role adaptation often leads to optimized coordination, improved task coverage, and enhanced robustness in complex and unpredictable scenarios [3, 4]. Nevertheless, the underlying difficulty in coordinating roles and decision-making across multiple agents in swarms poses significant challenges. For example, traditional control-based algorithms [5, 6] are contingent on fully connected topologies and lack the required adaptability and scalability to large-scale dynamic environments, thus degrading the practicality in the real world [7]. Meanwhile, despite the incorporation of inter-agent communications [8, 9] and cooperation [10, 11], it still remains inevitable for decentralized Multi-Agent Reinforcement Learning (MARL) to yield conflicting roles and decisions [12, 13]. Given the inherent semantic reasoning capabilities and pretrained experience of Large Language Models (LLMs) [14, 15], they offer a promising alternative to mitigate the critical issues of conflicting roles and inconsistent decision-making that plague purely MARL-based approaches.
1.1 Related Works
1.1.1 MARL-Based UAV Swarm Control
@@
Deep Reinforcement Learning (DRL)[16] has significantly enhanced agent adaptability in complex tasks. However, in Multi-Agent Systems (MAS), critical challenges, such as environmental non-stationarity, rapidly expanding state spaces, and difficulty in credit assignment, hinder the ability of traditional methods to learn effective cooperative policies. To address this problem, the Centralized Training with Decentralized Execution (CTDE) paradigm is introduced with exemplary algorithms like MADDPG[17] and MAPPO[18]. During training, it also involves techniques such as policy distillation[19] and imitation learning[20] to improve coordination in complex obstacle-laden environments. Nevertheless, these approaches struggle to assess an individual agent’s contribution, as they typically optimize a global value function while neglecting the importance of localized utility. Value decomposition methods (e.g., VDN[10], QMIX[21], QTRAN[22]) ameliorate this issue by decomposing the joint value function, thereby enabling analysis of individuals’ contributions to cooperative decision-making. Furthermore, more advanced attention mechanisms or neural communication protocols (e.g., TarMAC[11], IMANet[23], DAACMP[24]) can boost the effectiveness of filter messages. However, the direct communication of numerical vectors[11, 23, 24], which lacks interpretability and cannot convey task semantics, leads to information redundancy and bandwidth bottlenecks [8, 9, 25, 26], and greatly limits algorithm generalization. Although the leader–follower architecture[27] enables role differentiation, its static role assignments lack the flexibility to adapt to dynamic environmental conditions and varying formation sizes. Therefore, researchers resort to integrating hierarchical control with consensus inference[28, 29, 2], so as to simplify the inference interpretability and boost convergence speed. However, agents in these methods commonly remain homogeneous, and the policy networks learned via CTDE cannot inherently leverage the advantages of the natural heterogeneity in swarms. Therefore, building a scalable, efficient, dynamically adaptive, and interpretable heterogeneous multi-agent collaboration mechanism for UAV swarm control remains an open problem.
1.1.2 LLM-Based Multi-Agent Systems
@@
LLMs have demonstrated near-human performance[14] in complex reasoning and planning tasks, providing new impetus for environment understanding and decision explainability in UAV swarms[15]. Leveraging vast prior knowledge, LLMs can not only be used for single-agent path planning (e.g., CoNavGPT[30], RoCo[31]), but also facilitate high-level communication with low-level trajectory planning, significantly improving task efficiency, adaptability, and generalization[32, 33, 34]. Moreover, in MetaGPT[35], CAMEL[36] and ChatDev[37], LLMs can decompose complex missions into a number of subtasks handled collaboratively by different agents, thereby reducing “hallucinations” [38] and enhancing the ability to solve complex problems. More importantly, these LLM-driven Multi-Agent (LLM-MA) systems[39, 40, 41] often customize LLMs into specialized or personalized roles. Therefore, through multi-agent collaboration, they replicate human-like collective intelligence and further enhance overall situational understanding and decision explainability, making inter-agent interactions more meaningful. However, LLM-based decision-making still heavily relies on prior knowledge and lacks exploration, often getting stuck in local optima[42]. Additionally, these methods leverage individual memory and self-evolution mechanisms to optimize agents independently, while neglecting potential synergistic effects arising from multi-agent interactions. Furthermore, applying LLM-MA to build global consensus for DS-CEFC remains scarce, though some studies[41] start to focus on consensus negotiation among agents in a simplified scenario. On the other hand, some approaches[39] utilize diverse “personalities”, empowered by the LLM, for creative collaboration; nevertheless, the underlying fixed definitions of roles still struggle to accommodate dynamic task switching in DS-CEFC. In other words, enabling dynamic role adaptation with balanced exploration and exploitation remains a critical challenge for DS-CEFC.
1.1.3 Integration of MARL and LLM
@@
There has been some light shed on deeply integrating MARL and LLMs[43]. Prominently, semantic capabilities in LLMs can be leveraged as natural language interfaces to bridge human feedback and MAS [44], but the consensus inference therein often relies more on human supervision than on autonomous collaboration. Other studies distill LLM knowledge into smaller executable models or empower LLMs with human-in-the-loop feedback for policy generation, so as to accelerate MARL training and improve performance in complex environments [45, 46, 47]. Nevertheless, these methods often depend on offline human annotations and feedback, making them ill-suited for dynamic real-time changes; meanwhile, agent roles remain homogeneous, with only the number of agents being scaled up. Some recent work treats LLMs as the core of heterogeneous agents in MARL[14], assigning different “personalities” to agents. But these roles tend to be fixed and cannot adapt to environmental changes in DS-CEFC. Additionally, deploying LLMs in UAV swarms significantly strains computational resources and energy, limiting their practical use [48]. Thus, it is meaningful to calibrate superior means to integrate MARL and LLMs, thus better harnessing heterogeneous swarm intelligence for improved generalization and adaptability in DS-CEFC.
1.2 Contributions
@@
To accomplish consensus inference of multiple UAVs with heterogeneous roles for DS-CEFC, we propose a novel LLM-MARL-integrated framework called RALLY (Role-Adaptive LLM-Driven Yoked navigation). Built on our previous work[2], which unifies target selection, obstacle-avoidance navigation, and flight-control execution, RALLY serves as a significantly enhanced high-level consensus inference module towards establishing role-adaptive cooperative navigation. Specifically, RALLY comprises two core modules: an LLM-based two-stage semantic reasoning module and a Role-value Mixing Network (RMIX)-based credit-distribution mechanism. The integration of semantic reasoning in LLM and policy learning in MARL makes it qualified for coordinating roles and decision-making across swarms for DS-CEFC. The main contributions of this work can be summarized as follows:
@@
  • To meet the distributed deployment and parallel inference demands on edge devices, we propose a capacity-migration algorithm that significantly reduces runtime memory footprint. We validate RALLY in the Multi-Particle Environment (MPE)[49] and the Software-In-The-Loop (SITL) platform[50] based on Gazebo-ROS-PX4. Experimental results demonstrate that RALLY outperforms the latest MARL[2] and pure LLM-driven approaches[30, 51] in terms of task completion, convergence speed, generalization, and interpretability.
1.3 Paper Structure
The remainder of this paper is organized as follows. Section 2 introduces the system model and formulates the problem. Section 3 outlines the algorithmic framework of RALLY by elaborating on the RMIX role assignment mechanism and context-migration algorithm. Section 4 presents experimental results and discussion. Finally, Section 5 summarizes this work with possible research directions.
2 System Model and Problem Formulation
2.1 System Model
Beforehand, commonly used variables are summarized in Table 1.
Table 1: Mainly used notations in this paper.
@@ 2.1 System Model
$\bm{p}_{e}^{t}$, $\bm{v}_{e}^{t}$ Enemy’s position/velocity
@@
${g^{\prime}}_{i}^{t}$ Initial goal intention of agent $i$
@@
$\text{LLM}_{\text{HC}}$ Two-stage LLM-based consensus inference
@@
where $\bm{p}_{i}^{t}=[p_{x_{i}}^{t},p_{y_{i}}^{t}]$ and $\bm{v}_{i}^{t}=[v_{x_{i}}^{t},v_{y_{i}}^{t}]$ are the 2D position and velocity of agent $i$, respectively, while $\bm{p}_{e}^{t}=[p_{x_{e}}^{t},p_{y_{e}}^{t}]$, $\bm{v}_{e}^{t}=[\bm{v}_{x_{e}}^{t},\bm{v}_{y_{e}}^{t}]$ are those of enemy agent. Based on its local observation $\bm{o}_{i}^{t}$, agent $i$ first generates a target intention ${g}_{i}^{t}\in\mathcal{T}$ and role assignment $k_{i}^{t}\in\mathcal{K}=\{\texttt{Commander},\texttt{Coordinator},\texttt{Executor}\}$ (i.e., $\bm{a}_{i}^{t}=({g}_{i}^{t},k_{i}^{t})$), and then exchanges these proposals and observation (denoted as $\bm{M}_{\text{i}}^{t}=(\bm{a}_{i}^{t},\bm{o_{i}^{t}})$) within its neighborhood set $\mathcal{N}_{i}^{t}\in\mathcal{N}$, ultimately forming a collective consensus $g_{i}^{t}\in\mathcal{T}$ through semantic negotiation. Notably, different roles are associated with distinct decision-making strategies: Commander focuses on maximizing individual rewards through independent decision-making, tending to select points yielding the highest personal return; Coordinator balances team and individual gains, giving priority to the Commander’s choices whenever necessary; Executor primarily adheres to the Coordinator’s guidance, reverting to its own strategy if necessary for task reliability. As discussed later, the heterogeneous role assignment contributes to the consensus inference of large-scale swarms.
@@
As shown in Fig. 2, we further adopt a hierarchical policy structure. The mid-layer policy $\bm{\pi}_{M}$ guides the obstacle avoidance and formation $\mathcal{F}_{c}$, consisting of different numbers $c\in[C_{\min},C_{\max}]$ via MARL-based navigation, while the low-layer standard PID [53] $\bm{\pi}_{L}$ steers the flight control of UAV dynamics. In this paper, we assume the availability of $\bm{\pi}_{M}$ and $\bm{\pi}_{L}$ [2], and focus on the learning of the high-level policy $\bm{\pi}_{H}$ only, which is responsible for the inferred consensus $\mathcal{G}^{t}:=\{g_{i}^{t},\forall i\in\mathcal{N}\}$.
Figure 3: Flowchart of the RALLY algorithm. (a) Taking formation $\mathcal{F}_{4}$ as an example, UAVs 1, 3, 4, and 7 adaptively form the team. (b) Based on local observations $\bm{o}_{i}^{t}$, each agent selects target intention ${g’}_{i}^{t}$ and role $k_{i}$. Based on background prompt descriptions $\text{X}_{\text{task}}$, reasoning logic $\text{M}_{\text{CoT}}$, and local observation prompts $\text{Y}_{\text{init}}(\bm{o}_{i}^{t})$, UAVs generate target intentions via policy $\text{LLM}_{\text{init}}$, and then determine roles $k_{i}$ through a role selection policy $\text{RL}_{\text{HI}}$. (c) After neighborhood communications, $\text{LLM}_{\text{cons}}$ refines the goal to reach a final, converged consensus decision $g_{i}^{t}$ (e.g., Region A), incorporating individual role preferences. (d) The resulting formation $\mathcal{F}_{4}$ then collectively navigates toward Region A.
2.2 Problem Formulation
We define the DS-CEFC task as Decentralized Partially Observable Markov Decision Process (Dec-POMDP) extended to a multi-agent setting with heterogeneous roles and hierarchical semantics. The joint system state at time $t$ is